OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 72 and the last term is –102?

OpenStudy (anonymous):

Let n = 30, a = 72 and l = -102. Then using the formula for the arithmetic progression, \[S = n(a + l)/2 = 30[72 + (-102)]/2=-450.\]

OpenStudy (anonymous):

oh sorry, i mean it is for the sum of an AP. still the answer is the same.

OpenStudy (anonymous):

you are right hehe